Sharp Company manufactures a product for which the following standards have been set: Standard Quantity Standard Price StandaRequired: 1. For direct materials: a. Compute the actual cost per foot of materials for March. b. Compute the price varianceRequired: 1. For direct materials: a. Compute the actual cost per foot of materials for March. b. Compute the price varianceRequired: 1. For direct materials: a. Compute the actual cost per foot of materials for March. b. Compute the price variance
Answer #1
Materials Quantity variance
(AQ used - Std Qty allowed)*SR
(x - 3200*3)*5 =2100U
5x - 48000 = 2100U
x =(2100+48000)/5
10020
a) actual cost = 60120/10020
6 per foot answer
(
b) Materials price variance
(Actual rate - standard rate)*AQ purchased
(6 - 5)*10,020
10020 U answer
Spending variance = 2100U+10020U
12120 U answer
2a) labor spending varian 3,500 U
labor efficiency varian 1,050 U
labor price variance 2450 U
labor rate variance = ( Actual rate - standard rate)*Actual hours
(53,900- x*4900) = 2,450
x = (53900-2450)/4900
10.50
Standard hours 10.5 per hour
b) Standard hours =
labor efficiency variance
(Actual hours - standard hours allowed)*Standard rate
(4900 -x) *10.50 = 1,050
51,450 - 10.5x = 1050
x                       = (51450-1050)/10.5
4800
Standard hours allowed 4,800 hours
c) Standard hours allowed per unit of product
4800/3200
1.5
Standard hours allowed per unit of product 1.5 hours per unit